OpenStudy (anonymous):

Write the quadratic equation in factored form. Be sure to write the entire equation. x^2 - 5x - 24 = 0

OpenStudy (anonymous):

do you know the quadratic equation

OpenStudy (anonymous):

-b plus or minus \[\sqrt{b^2-4ac}/2a\]

OpenStudy (anonymous):

a would be 1 b would be 5 and c would be 24 just plug them in and simplify

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

it isnt working. @91004775 i need am equation to plug in

OpenStudy (anonymous):

\[ax^2+bx=c=0\] \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[x^2 - 5x - 24 = 0\] \[a=1.b=-5,c=-24\]

OpenStudy (anonymous):

or try \[x^2 - 5x - 24 = 0\] \[(x-8)(x+3)=0\]

OpenStudy (anonymous):

@ladiesman217 Can you help?

OpenStudy (anonymous):

you just have to plug them into the equation given above your answers have yo be 8 and 3

OpenStudy (anonymous):

@ladiesman217 Can you help?